OpenStudy (anonymous):

how does one determine the continuity of the function (sqrt of t)(1+t)?

OpenStudy (anonymous):

\(f(t)=(1+t)\sqrt{t}\), like this?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

this is a product of 2 functions. since both are continuous over their domains, their product will also be continuous so long as t stays within the domain of both functions. the only point that really doesn't have continuity is t = 0. this is because \(\lim_{t \rightarrow 0}(1+t)\sqrt{t}\) DNE (Does Not Exist). It's a one-sided limit, that is \(\lim_{t \rightarrow 0^+}(1+t)\sqrt{t}\) exists and is 0. therefore f(t) will be continuous for all t>0.

OpenStudy (anonymous):

does that make any sense?

OpenStudy (anonymous):

you there?

OpenStudy (anonymous):

yes. i am thinking about what u said

OpenStudy (anonymous):

okay, sorry to disturb

OpenStudy (anonymous):

it makes sense. thank u

OpenStudy (anonymous):

i'm around helping others but if you have any questions or comments, just post and i'll return to help you

OpenStudy (anonymous):

oops, nevermind. hope i helped.